A movie theater sells adult tickets at $5.50 a ticket and $3.50 for a childs ticket. How many of each ticket was sold if 21 tickets were bought for $83.50
Answer by Paul(988)

adults = x
childern = y
x+y=21
y=21-x (equation 1)
5.5x+3.5y=83.5 (sub in equation 1)
5.5x+3.5(21-x)=83.5
5.5x+73.5-3.5x=83.5
2x+73.5-73.5=83.5-73.5
2x=10(divide both sides by 2)
x=5
21-5 = 16
Therefore there were 5 adults and 16 childerns